OpenStudy (vjt):

Which of the following is the slope between the points (-2, 4) and (2, -2)? -3/2 -2/3 3/2 2/3

OpenStudy (gorv):

slope =(y2-y1)/(x2-x1)

OpenStudy (gorv):

that is =(-2-4)/(2-(-2))

OpenStudy (gorv):

=-6/4=-3/2